# Orbit-based structural decomposition and stellar population recovery for edge-on barred galaxies

### v1

@@ 3.3 Tagging the stellar orbits with ages and metallicities
After decomposing the galaxy, we separate the bar orbits and bulge/disk/halo orbits into two distinct $\lambda_{z}$$R$ phase spaces. Following Zhu et al. (2020) and Jin et al. (2024), we then use the Voronoi 2D binning method (Cappellari & Copin, 2003) to divide orbits in each $\lambda_{z}$$R$ phase space into different orbit bundles, with each bundle containing $\gtrsim 0.5\%$ of the total orbit weight. This creates 100-200 bundles in total, and Fig. 5 illustrates this procedure for Au-23-85-50.
@@
where $N_{b}$ is the total number of orbit bundles, and $f_{k}^{i}$ is the luminosity contributed by the orbit bundle $k$ in aperture $i$. By matching the model-fitted maps with the observed data, the age $t_{k}$ and metallicity $Z_{k}$ of each orbit bundle can be determined.
Similar to the kinematics, the goodness of model-fitted age and metallicity maps is evaluated by their chi-square difference from observations, which are expressed as
@@
From the age and metallicity maps, the youngest outer disk’s age and the metal-poorest halo’s metallicity can be directly predicted. However, ages and metallicities observed in the central regions are mixtures of different stellar populations, revealing the possible presence of intrinsically more metal-rich and older stellar populations than those observed from the 2D maps. Thus, we set the age and metallicity ranges of orbits as
@@
We adopt the bounded-variable least squares method (implemented by Python Scipy [^3] https://scipy.org/), which is an optimisation algorithm that solves linear least-squares problems with bounded parameters, to minimise $\chi_{\rm age}^{2}$ and $\chi_{\rm met}^{2}$ individually. By solving the bounded-variable least squares, the ages and metallicities for all orbits can be derived and are taken as the best-fitting stellar population model. The results of the entire best-fitting model, including surface brightness, kinematics, ages, and metallicities, are shown in Fig. 6.
@@ 4 Model-predicted structures and their true counterparts for an individual case
Figure 7: Spatial distributions of stellar mass for model-decomposed structures in Au-23-85-50 compared with their true counterparts. The top three rows display the true mass surface densities of Au-23 projected onto the internal $x$-$y$, $x$-$z$, and $y$-$z$ planes, while the bottom three rows show the model-predicted results for Au-23-85-50, with the central mass density normalised to unity. From left to right, the panels show the mass surface densities of the entire galaxy, the bar, the bulge, the disk, and the stellar halo. The mass fractions of all components are calculated within the $\rm 60\times 60\times 30\,arcsec^{3}$ volume ($\rm 1\,arcsec=0.2\,kpc$; aligned with the axis ranges in the panels), with their values shown in the text.
Figure 8: Spatial distributions of stellar ages for model-decomposed structures in Au-23-85-50 compared with their true counterparts. The top three rows display the stellar age distributions of Au-23 projected onto the internal $x$-$y$, $x$-$z$, and $y$-$z$ planes, while the bottom three rows show the model-predicted results for Au-23-85-50. From left to right, the panels show the stellar age distributions of the entire galaxy, the bar, the bulge, the disk, and the stellar halo. To highlight the age distributions, we only plot pixels brighter than $1\%$ of the galaxy’s brightness pixel for the bar and bulge, and $0.1\%$ for the disk and halo. The mass-weighted mean stellar ages of the entire galaxy and all components are calculated within the $\rm 60\times 60\times 30\,arcsec^{3}$ volume ($\rm 1\,arcsec=0.2\,kpc$; aligned with the axis ranges in the panels), with their values shown in the text.
@@
Figure 10: Comparison of the true and model-predicted mass density profiles of different structures. From left to right, the panels display normalised mass density profiles for Au-18, Au-23, and Au-28. The black, red, blue, and green lines denote the bar, bulge, disk, and stellar halo components, respectively. The solid curves represent the true mass density profiles, while the dashed lines denote model-predicted profiles derived from mock data sets with different viewing angles.
Figure 11: One-to-one comparison of the true and model-predicted mass fractions of decomposed structures. From left to right, the panels display mass fractions for Au-18, Au-23, and Au-28. The black, red, blue, and green markers denote the bar, bulge, disk, and stellar halo components, respectively. The triangles, squares, pentagons, and circles indicate mock data sets with viewing angles $(\theta_{\rm T},\varphi_{\rm T})=(85^{\circ},50^{\circ})$, $(85^{\circ},80^{\circ})$, $(89^{\circ},50^{\circ})$, and $(89^{\circ},80^{\circ})$, respectively. The coloured error bars in each panel represent the uncertainties for different components. The dashed lines represent the equality of model-predicted fractions $f_{\rm true}$ and true fractions $f_{\rm model}$, while the dotted lines are $\pm 0.1$ away from the dashed lines.
Figure 12: Comparison of the true and model-predicted stellar age profiles of different structures. From left to right, the panels display age profiles for Au-18, Au-23, and Au-28. The black, red, blue, and green lines denote the bar, bulge, disk, and stellar halo components, respectively. The solid lines represent the true mass-weighted age profiles, while the dashed lines denote the corresponding model-predicted profiles derived from mock data sets with different viewing angles.
@@
Figure 15: One-to-one comparison of the true and model-predicted mass-weighted mean stellar metallicities of decomposed structures. The dashed lines represent the equality of model-predicted metallicities $Z_{\rm model}$ and true metallicities $Z_{\rm true}$, while the dotted lines are $\rm\pm 0.5\,Z_{\odot}$ away from the dashed lines. The panels and markers are consistent with Fig. 11.
@@
Using particle information from the simulated galaxy Au-23, we derive the true mass distributions by assuming a constant mass-to-light ratio $M_{*}/L=2$ (identical to that used when creating the mock surface brightness in Sect. 2), whereas the model-predicted distributions are calculated from orbit weights and trajectories in the model of Au-23-85-50. By projecting the 3D mass distributions onto the internal $x$-$y$, $x$-$z$, and $y$-$z$ planes, the comparison of true and model-predicted distributions is demonstrated in Fig. 7. The model successfully reconstructs a BP/X-shaped bar, a spheroidal bulge, a thin disk, and a spatially diffuse stellar halo, closely matching the true structures in all projections. We calculate the mass fraction of each component within the $\rm 60\times 60\times 30\,arcsec^{3}$ volume ($\rm 1\,arcsec=0.2\,kpc$). For this model, the mass fractions of all components are recovered with absolute biases $\leq 0.05$.
Then, we calculate the 3D mass-weighted stellar age distributions for the simulated galaxy Au-23 and the model of Au-23-85-50, and project the distributions on the $x$-$y$, $x$-$z$, and $y$-$z$ planes, as shown in Fig. 8. The mass-weighted mean stellar ages of the entire galaxy and different components are calculated within the $\rm 60\times 60\times 30\,arcsec^{3}$ volume. The model reproduces a young disk ($t_{\rm model}=\rm 5.5\,Gyr$), an old bulge ($t_{\rm model}=\rm 9.4\,Gyr$), an old halo ($t_{\rm model}=\rm 10.2\,Gyr$), and an intermediate-age bar ($t_{\rm model}=\rm 6.5\,Gyr$). These results generally match the truth, with biases $t_{\rm model}-t_{\rm true}=-0.6$ and $\rm 0.6\,Gyr$ for the bar and the bulge, respectively.
Similar to the age distributions, the mass-weighted stellar metallicity distributions and mean stellar metallicities for Au-23-85-50 are derived and presented in Fig. 9. The model reproduces the most metal-rich bar ($Z_{\rm model}=3.4\,Z_{\odot}$), the most metal-poor halo ($Z_{\rm model}=0.8\,Z_{\odot}$), and the relatively metal-rich bulge and disk ($Z_{\rm model}=2.6$ and $2.1\,Z_{\odot}$, respectively). These values agree with the true metallicities for all components, with the biases $|Z_{\rm model}-Z_{\rm true}|\leq 0.3\,Z_{\odot}$.
@@ 5.1 Uncertainties of model-predicted properties
For model-predicted masses from each mock data set, we estimate their uncertainties following the same approach as in our previous paper (Jin et al., 2025). First, we fix the orbit libraries and thus also the generated potential in the best-fitting model and randomly perturb the mock kinematic maps 1000 times based on their error maps. Then, we calculate the orbit weights to derive 1000 new $\chi^{2}$ values. The fluctuation within the $\pm 1\sigma$ ($68\%$; $\Delta\chi^{2}_{\rm CL}$) region of these $\chi^{2}$ values is defined as the $1\sigma$ confidence level. The mass variations in models within the $1\sigma$ confidence level ($\chi^{2}-\chi^{2}_{\rm min}\leq\Delta\chi^{2}_{\rm CL}$) are treated as uncertainties.
Since the model-predicted stellar ages are derived from the stellar orbits, their uncertainties originate from variations in both mock kinematic maps and age maps. For each mock data set, we first fix the age maps and calculate the orbit ages for models within the $1\sigma$ confidence level. The variations of model-predicted stellar ages are defined as uncertainties induced by the kinematic maps $\Delta t_{\rm kin}$. Then, we randomly perturb the mock age maps 1000 times based on their error maps and recalculate the orbit ages with unchanged orbit libraries and weights. The standard deviations of ages from these 1000 perturbations are taken as the uncertainties induced by the age maps $\Delta t_{\rm age}$. We assume that these two factors independently contribute to the uncertainties. Thus, the overall uncertainties $\Delta t_{\rm all}$ can be derived using the error propagation formula ($\Delta t_{\rm all}^{2}=\Delta t_{\rm kin}^{2}+\Delta t_{\rm age}^{2}$). Similarly, for stellar metallicities, the overall uncertainties $\Delta Z_{\rm all}$ are calculated following the same procedure ($\Delta Z_{\rm all}^{2}=\Delta Z_{\rm kin}^{2}+\Delta Z_{\rm met}^{2}$).
@@ 5.2 Mass density profiles and mass fractions of different structures
The mass density profiles derived from all mock data sets are shown in the top panels of Fig. 10. The mass profiles of bars and bulges are compact, while those of disks and halos are extended. There are obvious ‘holes’ in the inner disks, which are distinct from the exponential profiles widely used in photometric decompositions. Across all cases, our models well recover the mass profiles in inner bars/bulges and outer disks/halos. However, some uncertainties emerge in outer bars/bulges and inner disks/halos, which are due to the low mass contributions at these radii.
To statistically quantify the model results across all mock data sets, we present a one-to-one comparison of the true and model-predicted mass fractions in Fig. 11. These mass fractions are calculated within the $\rm 60\times 60\times 30\,arcsec^{3}$ volume ($\rm 1\,arcsec=0.2\,kpc$). The halo fractions are recovered with absolute biases of $|f_{\rm model}-f_{\rm true}|\leq 0.03$ in all cases. For the bulge fractions, 10 of 12 cases show biases below 0.05, while the other two cases remain below 0.1. The absolute biases for the bar and disk fractions reach maximums of 0.15 and 0.14, respectively. On average, the models overestimate disk fractions by $f_{\rm model}-f_{\rm true}=0.06$ and underestimate bar fractions by $f_{\rm model}-f_{\rm true}=-0.06$, while bulge and halo fractions do not have significant biases. These discrepancies are mainly attributed to the model biases of the dynamical bar lengths, which are adopted to separate the bar/bulge and disk/halo components.
@@ 5.3 Stellar age profiles and mean stellar ages of different structures
Using the weights and ages assigned to the stellar orbits in the modelling, we calculate the mass-weighted stellar age profiles for different structures, as shown in Fig. 12. For the bars and disks, both the original and revised models generally reproduce their negative age gradients. The bars and disks exhibit a continuously negative age gradient, indicating the disk origin of the bar; this trend is generally matched by our model. For the bulges and halos, the predicted gradients have larger variations across different mock data sets, indicating higher uncertainties. However, the age difference between the bars and the bulges can still be identified: all three galaxies have bulges older than their bars at all radii, with the difference being the largest for Au-28 because its bar was recently formed ($\sim\rm 2.8\,Gyr$; Fragkoudi et al., 2025).
We calculate the mass-weighted mean stellar ages of the decomposed structures within the $\rm 60\times 60\times 30\,arcsec^{3}$ volume, and then compare them with the true values in Fig. 13. Although the original models slightly underestimate bar fractions and overestimate disk fractions, the mean stellar ages of different structures still align statistically with the true values. The mean age biases $t_{\rm model}-t_{\rm true}$ for the bar, bulge, disk, and halo components are 0.1-0.1, 0.4, 0.1-0.1, and $\rm 0.4\,Gyr$, respectively. We have relatively large uncertainties on the ages of the bulge and halo. One reason is that they are relatively minor components in their galaxies with lower mass fractions. Another reason is that they are older stellar populations, and we assigned larger age uncertainties to them when creating the mock maps. For all components, the absolute biases of the stellar ages are $\rm\lesssim 1\,Gyr$.

### v4

@@ 3.3 Tagging the stellar orbits with ages and metallicities
After decomposing the galaxy, we separated the bar orbits and bulge, disc, and halo orbits into two distinct $\lambda_{z}$$R$ phase spaces. Following Zhu et al. (2020) and Jin et al. (2024), we then used the Voronoi 2D binning method (Cappellari & Copin, 2003) to divide orbits in each $\lambda_{z}$$R$ phase space into different orbit bundles, with each bundle containing $\gtrsim 0.5\%$ of the total orbit weight. This created 100-200 bundles in total. Fig. 5 illustrates this procedure for Au-23-85-50.
@@
where $N_{b}$ is the total number of orbit bundles, and $f_{k}^{i}$ is the luminosity contributed by the orbit bundle, $k$, in aperture, $i$. By matching the model-fitted maps with the observed data, the age, $t_{k}$, and metallicity, $Z_{k}$, of each orbit bundle can be determined.
Similarly to the kinematics, the goodness of model-fitted age and metallicity maps is evaluated by their chi-square difference from observations, which are expressed as
@@
From the age and metallicity maps, the youngest outer disc’s age and the metal-poorest halo’s metallicity can be directly predicted. However, ages and metallicities observed in the central regions are mixtures of different stellar populations, revealing the possible presence of intrinsically more metal-rich and older stellar populations than those observed from the 2D maps. Thus, we set the age and metallicity ranges of orbits as
@@
We adopted the bounded-variable least squares method (implemented by Python Scipy [^3] https://scipy.org/), which is an optimisation algorithm that solves linear least-squares problems with bounded parameters, to minimise $\chi_{\rm age}^{2}$ and $\chi_{\rm met}^{2}$ individually. By solving the bounded-variable least squares, the ages and metallicities for all orbits can be derived and are taken as the best-fitting stellar population model. The results of the entire best-fitting model, including surface brightness, kinematics, ages, and metallicities, are shown in Fig. 6.
@@ 4 Model-predicted structures and their true counterparts for an individual case
Figure 7: Spatial distributions of stellar mass for model-decomposed structures in Au-23-85-50 compared with their true counterparts. The top three rows display the true mass surface densities of Au-23 projected onto the internal $x$-$y$, $x$-$z$, and $y$-$z$ planes, while the bottom three rows show the model-predicted results for Au-23-85-50, with the central mass density normalised to unity. From left to right, the panels show the mass surface densities of the entire galaxy, the bar, the bulge, the disc, and the stellar halo. The mass fractions of all components were calculated within the $\rm 60\times 60\times 30\,arcsec^{3}$ volume (aligned with the axis ranges in the panels), with their values shown in the text.
Figure 8: Spatial distributions of stellar ages for model-decomposed structures in Au-23-85-50 compared with their true counterparts. The top three rows display the stellar age distributions of Au-23 projected onto the internal $x$-$y$, $x$-$z$, and $y$-$z$ planes, while the bottom three rows show the model-predicted results for Au-23-85-50. From left to right, the panels show the stellar age distributions of the entire galaxy, the bar, the bulge, the disc, and the stellar halo. To highlight the age distributions, we only plot pixels brighter than $1\%$ of the galaxy’s brightness pixel for the bar and bulge, and $0.1\%$ for the disc and halo. The mass-weighted mean stellar ages of the entire galaxy and all components were calculated within the $\rm 60\times 60\times 30\,arcsec^{3}$ volume (aligned with the axis ranges in the panels), with their values shown in the text.
@@
Figure 10: Comparison of the true and model-predicted mass density profiles of different structures. From left to right, the panels display normalised mass density profiles for Au-18, Au-23, and Au-28. The black, red, blue, and green lines denote the bar, bulge, disc, and stellar halo components, respectively. The solid curves represent the true mass density profiles, while the dashed lines denote model-predicted profiles derived from mock datasets with different viewing angles.
Figure 11: One-to-one comparison of the true and model-predicted mass fractions of decomposed structures. From left to right, the panels display mass fractions for Au-18, Au-23, and Au-28. The black, red, blue, and green markers denote the bar, bulge, disc, and stellar halo components, respectively. The triangles, squares, pentagons, and circles indicate mock datasets with viewing angles $(\theta_{\rm T},\varphi_{\rm T})=(85^{\circ},50^{\circ})$, $(85^{\circ},80^{\circ})$, $(89^{\circ},50^{\circ})$, and $(89^{\circ},80^{\circ})$, respectively. The coloured error bars in each panel represent the uncertainties for different components. The dashed lines represent the equality of model-predicted fractions $f_{\rm true}$ and true fractions $f_{\rm model}$, while the dotted lines are $\pm 0.1$ away from the dashed lines.
Figure 12: Comparison of the true and model-predicted stellar age profiles of different structures. From left to right, the panels display age profiles for Au-18, Au-23, and Au-28. The black, red, blue, and green lines denote the bar, bulge, disc, and stellar halo components, respectively. The solid lines represent the true mass-weighted age profiles, while the dashed lines denote the corresponding model-predicted profiles derived from mock datasets with different viewing angles.
@@
Figure 15: One-to-one comparison of the true and model-predicted mass-weighted mean stellar metallicities of decomposed structures. The dashed lines represent the equality of model-predicted metallicities, $Z_{\rm model}$, and true metallicities, $Z_{\rm true}$, while the dotted lines are $\rm\pm 0.5\,Z_{\odot}$ away from the dashed lines. The panels and markers are consistent with Fig. 11.
@@
Using particle information from the simulated galaxy Au-23, we derived the true mass distributions by assuming a constant mass-to-light ratio $M_{*}/L=2$ (identical to that used when creating the mock surface brightness in Sect. 2), whereas the model-predicted distributions were calculated from orbit weights and trajectories in the model of Au-23-85-50. By projecting the 3D mass distributions onto the internal $x$-$y$, $x$-$z$, and $y$-$z$ planes, the comparison of true and model-predicted distributions is demonstrated in Fig. 7. The model successfully reconstructed a BP/X-shaped bar, a spheroidal bulge, a thin disc, and a spatially diffuse stellar halo, closely matching the true structures in all projections. We calculated the mass fraction of each component within the $\rm 60\times 60\times 30\,arcsec^{3}$ volume. For this model, the mass fractions of all components are recovered with absolute biases $\leq 0.05$.
Then, we calculated the 3D mass-weighted stellar age distributions for the simulated galaxy Au-23 and the model of Au-23-85-50. We projected the distributions on the $x$-$y$, $x$-$z$, and $y$-$z$ planes, as shown in Fig. 8. The mass-weighted mean stellar ages of the entire galaxy and different components were calculated within the $\rm 60\times 60\times 30\,arcsec^{3}$ volume. The model reproduced a young disc ($t_{\rm model}=\rm 5.5\,Gyr$), an old bulge ($t_{\rm model}=\rm 9.4\,Gyr$), an old halo ($t_{\rm model}=\rm 10.2\,Gyr$), and an intermediate-age bar ($t_{\rm model}=\rm 6.5\,Gyr$). These results generally match the truth, with biases $t_{\rm model}-t_{\rm true}=-0.6$ and $\rm 0.6\,Gyr$ for the bar and the bulge, respectively.
Similarly to the age distributions, the mass-weighted stellar metallicity distributions and mean stellar metallicities for Au-23-85-50 are presented in Fig. 9. The model reproduced the most metal-rich bar ($Z_{\rm model}=3.4\,Z_{\odot}$), the most metal-poor halo ($Z_{\rm model}=0.8\,Z_{\odot}$), and the relatively metal-rich bulge and disc ($Z_{\rm model}=2.6$ and $2.1\,Z_{\odot}$, respectively). These values agree with the true metallicities for all components, with all components showing biases, expressed as $|Z_{\rm model}-Z_{\rm true}|\leq 0.3\,Z_{\odot}$.
@@ 5.1 Uncertainties of model-predicted properties
For model-predicted masses from each mock dataset, we estimated their uncertainties following the same approach as in our previous paper (Jin et al., 2025). First, we fixed the orbit libraries and thus also the generated potential in the best-fitting model and randomly perturbed the mock kinematic maps 1000 times based on their error maps. Then, we calculated the orbit weights to derive 1000 new $\chi^{2}$ values. The fluctuation within the $\pm 1\sigma$ ($68\%$; $\Delta\chi^{2}_{\rm CL}$) region of these $\chi^{2}$ values is defined as the $1\sigma$ confidence level. The mass variations in models within the $1\sigma$ confidence level ($\chi^{2}-\chi^{2}_{\rm min}\leq\Delta\chi^{2}_{\rm CL}$) are treated as uncertainties.
Since the model-predicted stellar ages are derived from the stellar orbits, their uncertainties originate from variations in both mock kinematic maps and age maps. For each mock dataset, we first fixed the age maps and calculated the orbit ages for models within the $1\sigma$ confidence level. The variations among the model-predicted stellar ages are defined as uncertainties induced by the kinematic maps, $\Delta t_{\rm kin}$. Then, we randomly perturbed the mock age maps 1000 times based on their error maps and recalculated the orbit ages with unchanged orbit libraries and weights. The standard deviations of ages from these 1000 perturbations are taken as the uncertainties induced by the age maps, $\Delta t_{\rm age}$. We assumed that these two factors independently contribute to the uncertainties. Thus, the overall uncertainties, $\Delta t_{\rm all}$, could be derived using the error propagation formula ($\Delta t_{\rm all}^{2}=\Delta t_{\rm kin}^{2}+\Delta t_{\rm age}^{2}$). Similarly, for stellar metallicities, the overall uncertainties, $\Delta Z_{\rm all}$, were calculated following the same procedure ($\Delta Z_{\rm all}^{2}=\Delta Z_{\rm kin}^{2}+\Delta Z_{\rm met}^{2}$).
@@ 5.2 Mass density profiles and mass fractions of different structures
The mass density profiles derived from all mock datasets are shown in the top panels of Fig. 10. The mass profiles of bars and bulges are compact, while those of discs and halos are extended. There are obvious ‘holes’ in the inner disc, which are distinct from the exponential profiles widely used in photometric decompositions. Across all cases, our models aptly recovered the mass profiles in inner bars, inner bulges, outer discs, and outer halos. However, some uncertainties emerge in outer bars, outer bulges, inner discs, and inner halos, which are due to the low mass contributions at these radii.
To statistically quantify the model results across all mock datasets, we present a one-to-one comparison of the true and model-predicted mass fractions in Fig. 11. These mass fractions have been calculated within the $\rm 60\times 60\times 30\,arcsec^{3}$ volume. The halo fractions are recovered with absolute biases of $|f_{\rm model}-f_{\rm true}|\leq 0.03$ in all cases. For the bulge fractions, 10 out of 12 cases show biases below 0.05, while the other two cases remain below 0.1. The absolute biases for the bar and disc fractions reach maximums of 0.15 and 0.14, respectively. On average, the models overestimate the disc fractions by $f_{\rm model}-f_{\rm true}=0.06$ and underestimate the bar fractions by $f_{\rm model}-f_{\rm true}=-0.06$, while the bulge and halo fractions do not exhibit significant biases. These discrepancies are mainly attributed to the model biases of the dynamical bar lengths, which have been adopted to separate the bar+bulge and disc+halo components.
@@ 5.3 Stellar age profiles and mean stellar ages of different structures
Using the weights and ages assigned to the stellar orbits in the modelling, we calculated the mass-weighted stellar age profiles for different structures, as shown in Fig. 12. For bars and discs, the models generally reproduced their negative age gradients. The bars and discs exhibit a consistently negative age gradient, indicating the disc origin of the bar; this trend is generally matched by our model. For bulges and halos, the predicted gradients have larger variations across different mock datasets, indicating higher uncertainties. However, the age difference between the bars and the bulges can still be identified: all three galaxies have bulges older than their bars at all radii, with the difference being the largest for Au-28 because its bar was recently formed ($\sim\rm 2.8\,Gyr$; Fragkoudi et al., 2025).
We calculated the mass-weighted mean stellar ages of the decomposed structures within the $\rm 60\times 60\times 30\,arcsec^{3}$ volume and then compared them with the true values in Fig. 13. Although the models slightly underestimate the bar fractions and overestimate disc fractions, the mean stellar ages of different structures are still statistically aligned with the true values. The mean age biases $t_{\rm model}-t_{\rm true}$ for the bar, bulge, disc, and halo components are 0.1-0.1, 0.4, 0.1-0.1, and $\rm 0.4\,Gyr$, respectively. We have relatively large uncertainties on the ages of the bulge and halo. One reason is that they are relatively minor components in their galaxies with lower mass fractions. Another reason is that they are older stellar populations and we assigned larger age uncertainties to them when creating the mock maps. For all components, the absolute biases of the stellar ages are $\rm\lesssim 1\,Gyr$.